OpenStudy (anonymous):

An empty rectangular water tank has a flat rectangular base that measures 2 feet by 4 feet, and a height of 5 feet. If 10 cubic feet of water is poured into the tank, about how many feet will the water's surface be above the tank's base? less than a foot between 1 and 2 feet between 2 and 3 feet between 3 and 4 feet between 4 and 5 feet

OpenStudy (ghazi):

volume of rectangular tank= length* width* height= 2*4*5=40 cubic feet now volume of water stored in it= 10 cubic feet so remaining volume= 30 cubic feet now to find the height of 10 cubic feet use V= l*w*h 10= 2*4*h h=10/8

OpenStudy (anonymous):

between 1 and 2 feet?

OpenStudy (ghazi):

yes